A rectangular piece of land measures 0.7 km by 0.5 km. Each side is to be fenced with 4 rows of wires. What is the length of the wire needed?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Perimeter of the field = 2 (Length + Breadth)
= 2 (0.7 + 0.5) = 2 (1.2)
= 2 × 1.2
= 2.4 km
Each side is to be fenced with 4 rows = 4 × 2.4
= 9.6 km
∴Total length of the required wire is 9.6 km

Thanks